40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 17 18:08:45 2017, Dhruva, Update, Optical Levers, Beam Profiling Setup 1.jpg2.jpg3.jpg4.jpg5.jpg
    Reply  Mon May 22 10:53:02 2017, Dhruva, Update, Optical Levers, Beam Profiling Results  6x
       Reply  Tue May 23 10:27:24 2017, Dhruva, Update, Optical Levers, Beam Profiling Results  6x
          Reply  Tue May 23 15:22:04 2017, rana, Update, Optical Levers, Beam Profiling Results  
             Reply  Tue May 30 18:31:54 2017, Dhruva, Update, Optical Levers, Beam Profiling Results  plots.pdfspot_size_y.pdf
                Reply  Thu Jun 8 12:43:42 2017, Dhruva, Update, Optical Levers, Beam Profiling Results  profile.pdf
          Reply  Tue May 23 16:33:00 2017, Steve, Update, Optical Levers, Beam Profiling Results  
Message ID: 12997     Entry time: Wed May 17 18:08:45 2017     Reply to this: 13002
Author: Dhruva 
Type: Update 
Category: Optical Levers 
Subject: Beam Profiling Setup 

Andrew and I set up the razor blade beam profiling experiment for He-Ne lasers on the "SP" table.  Once I receive the laser safety training, I will make power measurements and fit it to an erfc curve from which I will calculate the gaussian profile of the beam. I'm attaching some pictures of the setup. 

Least count of the micrometer - 2 microns 

Laser  : Lumentum 22037130:1103P

Photodetector : Thor Labs PDA100A

Attachment 1: 1.jpg  2.478 MB  Uploaded Wed May 17 19:20:18 2017  | Hide | Hide all
1.jpg
Attachment 2: 2.jpg  2.648 MB  Uploaded Wed May 17 19:20:18 2017  | Hide | Hide all
2.jpg
Attachment 3: 3.jpg  2.763 MB  Uploaded Wed May 17 19:20:18 2017  | Hide | Hide all
3.jpg
Attachment 4: 4.jpg  2.567 MB  Uploaded Wed May 17 19:20:18 2017  | Hide | Hide all
4.jpg
Attachment 5: 5.jpg  2.674 MB  Uploaded Wed May 17 19:20:18 2017  | Hide | Hide all
5.jpg
ELOG V3.1.3-